# Unsupervised Improved MVDR Beamforming for Sound Enhancement

### v2

@@ Abstract
Neural networks have recently become the dominant approach to sound separation. Their good performance relies on large datasets of isolated recordings. For speech and music, isolated single channel data are readily available; however the same does not hold in the multi-channel case, and with most other sound classes. Multi-channel methods have the potential to outperform single channel approaches as they can exploit both spatial and spectral features, but the lack of training data remains a challenge. We propose unsupervised improved minimum variation distortionless response (UIMVDR), which enables multi-channel separation to leverage in-the-wild single-channel data through unsupervised training and beamforming. Results show that UIMVDR generalizes well and improves separation performance compared to supervised models, particularly in cases with limited supervised data. By using data available online, it also reduces the effort required to gather data for multi-channel approaches.
@@ 1 Introduction
In noisy and reverberant environments, SED [7], SSL [8, 9], SC [10] and SR [11] performance degrades because the target signal is mixed with interfering signals. This is especially challenging for robots because their actuators may generate noise and they may interact in noisy and reverberant indoor environments. To alleviate this, sound separation or sound enhacement can be used prior to recognition to estimate isolated sounds for use in SED, SSL, SC and SR. Recently, deep learning algorithms have achieved greatly improved sound source separation performance, leading to improvement in downstream recognition tasks such as SED [12], SSL [8], SC [10] and SR [11].
@@
Multi-channel approaches often use datasets that are synthetically created using single channel data and room impulse responses (RIRs), in order to provide isolated ground truth sources for supervised training. However, the generated data can differ from real data because of the challenges associated with accurately simulating the reverberation characteristics of an actual room. Custom datasets for a specific microphone array can also be created but this is time consuming, and models trained on such data generalize poorly to other microphone arrays. The use of beamformers using signal estimates from single channel approaches can overcome these challenges [14, 15]. Initially, single channel approaches can extract a mask of the signal of interest using spectral features with a reference channel. This mask can then be applied to each channel of the multi-channel input to obtain an estimate of the signal of interest. Subsequently, this estimated signal can be refined using Minimum Variance Distortionless Response (MVDR) beamforming [16], which uses the spatial features to enhance the target signal [14, 15, 17]. The deep learning algorithm used in this approach can be trained on single-channel input and therefore eliminates the need of multi-channel datasets for processing input from multiple microphones.
@@
Unsupervised Improved MVDR beamforming combines the multi-channel beamforming approach with the single-channel unsupervised approach to enhance a sound of interest. This enables multi-channel sound enhancement to benefit from large real single channel databases like recordings made on phones or recordings available on video sharing services. We also propose a new dataset to evaluate supervised, unsupervised, single and multi-channel sound separation algorithms. The code and evaluation dataset are available online  [^1] https://github.com/introlab/uimvdr.
@@ 2 Proposed Method
The proposed method is shown in figure 1. The mixture is first encoded using a Short-Term Fourier Transform (STFT). The STFT is used as the encoder for compatibility with the MVDR beamformer and because it performs better for sound separation as seen in Kavelerov et al. [22]. They also report that, for windows longer than 5 ms, STFT outperforms learnt encoders. In order to perform well, beamforming needs a window size that is on the order of the reverberation length [23]. Window sizes often used are 64 [24] and 32 ms [25]. STFT encoders also tend to perform better in reverberant environments [26]. A single-channel unsupervised neural network is then used to separate the sources. Beamforming is applied using the estimated target signal to improve the enhancement. Finally, the inverse Short-Term Fourier Transform (ISTFT) is used to decode the enhanced signal and obtain the final prediction in the time domain. This technique can also be used for source separation by beamforming on every output.
@@ 2.2 Efficient MixIT
Separation models can be trained using full supervision [27] but this remains difficult for general purpose sound enhancement and sound separation as the clean isolated sources and the mixture are rarely available in real recordings, without significant cross talk. Unsupervised training enables us to use noisy recordings made with everyday devices. This increases the amount of data available for training and eases data collection. To train the neural network without supervision, we used the unsupervised framework Mixture Invariant Training (MixIT) [19, 20]. Unsupervised training also helps with generalizing to multiple environments [19]. MixIT combines two or more mixtures to create a mixture of mixtures (MoM). The MoM is fed as the input of the separation model. The model then predicts the separated sources. Using a mixing matrix $A$, the separated sources are assigned to one of the original mixtures. This mixing matrix is obtained by computing the loss on every reconstructed mixtures and the original mixtures. The mixture with the best loss is then selected for every prediction. Although this works well, it can be quite computationally expensive to calculate the loss for every possible assignment. To address this, [20] proposes an efficient version of MixIT using the least-squares algorithm:
@@ 2.3 MVDR Beamforming
While using a weakly supervised deep learning model achieves good results for sound enhancement and separation, there is room for improvement. In fact, the mask prediction can sometimes omit part of the target signal or have residual noise. To address this, we use the predicted signals to compute spatial covariance matrices (SCM) (${\bf{\Phi}_{\mathbf{\hat{N}\hat{N}}}}$, ${\bf{\Phi}_{\mathbf{\hat{X}\hat{X}}}}$) and then beamform in that direction using MVDR, as follows:
@@
where {}$\{...\}^{H}$ stands for the Hermitian operator and $\mathbf{u}$ is a one-hot vector indicating the reference microphone. This is possible when multiple channels are available, assuming the sources come from different directions. MVDR beamforming also ensures the linearity of the estimate which is useful when enhancement is applied upstream of another algorithm. To further improve beamforming estimates, a minimum floor post-masking is used [14].
@@ 3 Experiments and discussion
Tables 1 and 2 present the results of different training methods on 4 test datasets. The 3 from MCFSTD (ReSpeaker, Kinect and 16Sounds) and the final one is the test dataset from FSD50K or Librispeech (Single) [33]. To create the Single test dataset and supervised training datasets, isolated targets are necessary. For target signals in bark enhancement, we use audio samples in FSD50K that only has the bark label as target. For speech, we use samples from Librispeech augmented with RIRs from BIRD [34]. For the non-target signals in both cases, we use the samples in FSD50K that does not include the target class. To create MoMs, we iterate 200 times on the Single bark target mixtures, mixing them with different interference. This is done 10 times for every other datasets. We do not iterate for target only results. For testing and training, the MoM contains 2 to 4 mixtures. The gain of every mixture is normalized randomly to between -5 and 5 dB to add more robustness in the network. In both speech and bark enhancement, there is a noticeable improvement across all cases when beamforming is applied, as opposed to relying solely on the network predictions. The weighting leads to slight improvements in some cases.
@@
The robustness in change of domain with unsupervised training compared to supervised training can also be observed in the results as first noted in [19]. It is possible to observe this by subtracting the SI-SDRi of the in-domain test dataset (Single) with the SI-SDRi of the out-of-domain datasets (MCFSTD). On average, we note a larger performance drop of 0.31 dB for supervised training in contrast to unsupervised training. However, this is not possible to observe for models trained on AudioSet as the Single dataset is also out-of-domain. This is crucial for training a neural network that will maintain high performance once deployed in real-world scenarios. Particularly when using training data from the exact target domain is impossible.
When looking at the results for speech enhancement it is also possible to observe that the supervised network outperforms the unsupervised networks trained on AudioSet [35]. However, once the beamforming is applied, the inverse is observed in the SI-SDRi for the ReSpeaker and 16Sounds and in the PESQ for the MCFSTD datasets. We hypothesize that this is because the unsupervised networks seems to be less aggressive in suppressing interference than the supervised networks. This results in a reduced presence of the target in the noise SCM but an increased amount of interference in the target SCM compared to the supervised network prediction. This seems to help the MVDR beamformer as its objective is to minimize the power of the noise while constraining the distortion in the target direction [16]. Having a smoother noise SCM can also contribute to a better numerical stability when computing the inverse of the noise SCM. This is important when using a mask predicted by a neural network because the SCM can be very sparse in some frequencies, as first noted in [15].

### v4

@@ Abstract
Neural networks have recently become the dominant approach to sound separation. Their good performance relies on large datasets of isolated recordings. For speech and music, isolated single channel data are readily available; however, the same does not hold in the multi-channel case, and with most other sound classes. Multi-channel methods have the potential to outperform single channel approaches as they can exploit both spatial and spectral features, but the lack of training data remains a challenge. We propose unsupervised improved minimum variation distortionless response (UIMVDR), which enables multi-channel separation to leverage in-the-wild single-channel data through unsupervised training and beamforming. Results show that UIMVDR generalizes well and improves separation performance compared to supervised models, particularly in cases with limited supervised data. By using data available online, it also reduces the effort required to gather data for multi-channel approaches.
@@ 1 Introduction
In noisy and reverberant environments, SED [7], SSL [8, 9], SC [10] and SR [11] performance degrades because the target signal is mixed with interfering signals. This is especially challenging for robots because their actuators may generate noise and they may interact in noisy and reverberant indoor environments. To alleviate this, sound separation or sound enhancement can be used prior to recognition to estimate isolated sounds for use in SED, SSL, SC and SR. Recently, deep learning algorithms have achieved greatly improved sound source separation performance, leading to improvement in downstream recognition tasks such as SED [12], SSL [8], SC [10] and SR [11].
@@
Multi-channel approaches often use datasets that are synthetically created using single channel data and room impulse responses (RIRs), in order to provide isolated ground truth sources for supervised training. However, the generated data can differ from real data because of the challenges associated with accurately simulating the reverberation characteristics of an actual room. Custom datasets for a specific microphone array can also be created, but this is time consuming, and models trained on such data generalize poorly to other microphone arrays. The use of beamformers using signal estimates from single channel approaches can overcome these challenges [14, 15]. Initially, single channel approaches can extract a mask of the signal of interest using spectral features with a reference channel. This mask can then be applied to each channel of the multi-channel input to obtain an estimate of the signal of interest. Subsequently, this estimated signal can be refined using minimum variance distortionless response (MVDR) beamforming [16], which uses the spatial features to enhance the target signal [14, 15, 17]. The deep learning algorithm used in this approach can be trained on single-channel input and therefore eliminates the need of multi-channel datasets for processing input from multiple microphones.
@@
Unsupervised improved MVDR beamforming (UIMVDR) combines the multi-channel beamforming approach with the single-channel unsupervised approach to enhance a sound of interest. This enables multi-channel sound enhancement to benefit from large real single channel databases like recordings made on phones or recordings available on video sharing services. We also propose a new dataset to evaluate supervised, unsupervised, single-channel, and multi-channel sound separation algorithms. The code and evaluation dataset are available online  [^1] https://github.com/introlab/uimvdr.
@@ 2 Proposed Method
The proposed method is shown in figure 1. The mixture is first encoded using a short-time Fourier transform (STFT). The STFT is used as the encoder for compatibility with the MVDR beamformer and because it performs well for sound separation as seen in Kavelerov et al. [22]. They also report that, for windows longer than 5 ms, STFT outperforms learned encoders. In order to perform well, beamforming needs a window size that is on the order of the reverberation length [23]. Window sizes often used are 64 [24] and 32 ms [25]. STFT encoders also tend to perform better in reverberant environments [26]. A single-channel unsupervised neural network is then used to separate the sources. Beamforming is applied using the estimated target signal to improve the enhancement. Finally, the inverse short-time Fourier Transform (ISTFT) is used to decode the enhanced signal and obtain the final prediction in the time domain. This technique can also be used for multi-source separation by using each estimated source as a beamforming target.
@@ 2.2 Efficient MixIT
Separation models can be trained using full supervision [27] but this remains difficult for general purpose sound enhancement and sound separation as the clean isolated sources and the mixture are rarely available in real recordings, without significant cross talk. Unsupervised training enables us to use noisy recordings made with everyday devices. This increases the amount of data available for training and eases data collection. To train the neural network without supervision, we used the unsupervised framework, mixture invariant training (MixIT) [19, 20]. Unsupervised training also helps with generalizing to multiple environments [19]. MixIT combines two or more mixtures to create a mixture of mixtures (MoM). The MoM is fed as the input of the separation model. The model then predicts the separated sources. Using a mixing matrix $A$, the separated sources are assigned to one of the original mixtures. For each possible mixing matrix, the assigned sources are summed to form reconstructed mixtures. The mixing matrix is selected to minimize the loss between the reconstructed mixtures and the original mixtures. Although this works well, it can be quite computationally expensive to calculate the loss for every possible assignment. To address this, [20] proposes an efficient version of MixIT using the least-squares algorithm:
@@ 2.3 MVDR Beamforming
While using a weakly supervised deep learning model achieves good results for sound enhancement and separation, there is room for improvement. In fact, the mask prediction can sometimes omit part of the target signal or produce residual noise. To address this, we use the predicted signals to compute spatial covariance matrices (SCMs) (${\bf{\Phi}_{\mathbf{\hat{N}\hat{N}}}}$, ${\bf{\Phi}_{\mathbf{\hat{X}\hat{X}}}}$) and then beamform in that direction using MVDR, as follows:
@@
where {}$\{...\}^{H}$ stands for the Hermitian operator and $\mathbf{u}$ is a one-hot vector indicating the reference microphone. This is possible when multiple channels are available, assuming the sources come from different directions. MVDR beamforming also ensures the linearity of the estimate which is useful when enhancement is applied upstream of another algorithm. To further improve beamforming estimates, a minimum floor post-masking filter is used [14].
@@ 3 Experiments and discussion
Tables 1 and 2 present the results of different training methods on 4 test datasets. The 3 from MCFSTD (ReSpeaker, Kinect and 16Sounds) and the final one is the test dataset from FSD50K or Librispeech (Single) [33]. To create the Single test dataset and supervised training datasets, isolated targets are necessary. For target signals in bark enhancement, we use audio samples in FSD50K that only has the bark label as target. For speech, we use samples from Librispeech augmented with RIRs from BIRD [34]. For the non-target signals in both cases, we use the samples in FSD50K that does not include the target class. To create MoMs, we iterate 200 times on the Single bark target mixtures, mixing them with different interference. This is done 10 times for each other dataset. We do not iterate for target only results. For testing and training, the MoM contains 2 to 4 mixtures. The gain of every mixture is normalized randomly to between -5 and 5 dB to add more robustness in the network. In both speech and bark enhancement, there is a noticeable improvement across all cases when beamforming is applied, as opposed to relying solely on the network predictions. The weighting leads to slight improvements in some cases.
@@
The robustness in change of domain with unsupervised training compared to supervised training can also be observed in the results as first noted in [19]. It is possible to observe this by subtracting the SI-SDRi of the in-domain test dataset (Single) with the SI-SDRi of the out-of-domain datasets (MCFSTD). On average, we note a larger performance drop of 0.31 dB for supervised training in contrast to unsupervised training. However, this is not possible to observe for models trained on AudioSet as the Single dataset is also out-of-domain. Domain robustness is crucial for training a neural network that will maintain high performance once deployed in real-world scenarios, particularly when using training data from the exact target domain is impossible.
When looking at the results for speech enhancement, it is also possible to observe that the supervised network outperforms the unsupervised networks trained on AudioSet [35]. However, once the beamforming is applied, the inverse is observed in the SI-SDRi for the ReSpeaker and 16Sounds and in the PESQ for the MCFSTD datasets. We hypothesize that this is because the unsupervised networks seems to be less aggressive in suppressing interference than the supervised networks. This results in a reduced presence of the target in the noise SCM but an increased amount of interference in the target SCM compared to the supervised network prediction. This seems to help the MVDR beamformer as its objective is to minimize the power of the noise while constraining the distortion in the target direction [16]. Having a smoother noise SCM can also contribute to a better numerical stability when computing the inverse of the noise SCM. This is important when using a mask predicted by a neural network because the SCM can be very sparse in some frequencies, as first noted in [15].